In an economy, the working-age population is 400 million. Of this total, 320 million workers are employed. 12 million workers are unemployed. 56 million workers are not available for work (homemakers, full-time students, etc.). 8 million workers are available for work but are discouraged and thus are not seeking work. 4 million workers are available for work but are not currently seeking work due to transportation or childcare problems The unemployment rate in this economy is | %. (Round your answer to one decimal place.)
Answer #1

Unemployment Rate = (Number of unemployed / Labour force) × 100

Labour force = number of unemployed people + number of employed people.

Labour force = 320+12 = 332 million

Unemployment Rate = (12/332)×100 = 3.6%